# On modules whose dual is of finite Gorenstein dimension: The $k$-torsion, homological criteria, and applications in modules of differentials of order $n$

### v4

@@ 3. A generalization of a theorem of Evans and Griffith
Let $R$ be a local ring of depth $t$, $M$ be a non-zero $R$-module and $0\leq k\leq t$ be an integer. If $M$ is a $k$-syzygy of an $R$-module of depth 00, then $\operatorname{depth}_{R}(M)=k$.
@@
It is sufficient to consider $k\geq 1$. We proceed by induction on $k$. Suppose $k=1$. Then $t\geq 1$ and there exists an exact sequence $0\to M\to F_{0}\to L\to 0,$ where $F_{0}$ is free $R$-module and $L$ has of depth 00. The depth lemma gives us the inequalities
@@
Since $t\geq 1$, it follows from (3.1) that $\operatorname{depth}_{R}(M)\leq 1$. Thus (3.2) say us that $\operatorname{depth}_{R}(M)=1$.
Now suppose $1<k\leq t$. Since $M$ is a $k$-syzygy, there exists an exact sequence
@@
where $\operatorname{depth}_{R}(L)=0$. Then it induces exact sequences
@@
(3.4) $0\to N\to F_{k-2}\to\cdots\to F_{0}\to L\to 0$
If $N=0$, then (3.4) says that $\operatorname{pd}_{R}(L)\leq k-2$ and, by the Auslander-Buchsbaum formula, $t\leq k-2,$ which does not occurs by assumption. So, $N\not=0$. Then by induction, we have that $\operatorname{depth}_{R}(N)=k-1$. Applying the depth formula in (3.3) we get that
@@
Thus, if $t\leq\operatorname{depth}_{R}(M)-1$, then $k-1\geq t$, which contradicts our assumption. Therefore, we must have $t>\operatorname{depth}_{R}(M)-1$. Consequently, $\operatorname{depth}_{R}(M)-1\leq t$, and thus $\operatorname{depth}_{R}(M)-1\leq k-1$, which implies $\operatorname{depth}_{R}(M)\leq k$. Therefore, $k=\operatorname{depth}_{R}(M)$. ∎
Below, we provide a characterization of a module to be $k$-torsionfree and $k$-syzygy, involving the condition $(\widetilde{S}_{k})$. This characterization generalizes a celebrated result by Evans and Griffith ([15, Theorem 3.8]).
@@
Let $R$ be a ring and let $k$ be a non-negative integer. Let $M$ be a non-zero $R$-module such that $M^{\ast}$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(R)$. Then the following conditions are equivalent:
@@
The implications $(1)\Rightarrow(2)\Rightarrow(3)$ follow from [50, Proposition 38]. Now let us prove $(3)\Rightarrow(1)$. We prove this by induction on $k$. If $k=0$, there is nothing to prove. Assume that $k\geq 1$. By induction, $M$ is $(k-1)$-torsionfree, that is $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $11\leq i\leq k-$ 1, and so it remains to conclude that $\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R)=0$.
By contradiction, assume that $\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R)\not=0$. Let $\mathfrak{p}\in\operatorname{Ass}_{R}(\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R))$. Then $\mathfrak{p}R_{\mathfrak{p}}\in\operatorname{Ass}_{R_{\mathfrak{p}}}\left(\operatorname{Ext}_{R_{\mathfrak{p}}}^{k}\left(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}}\right)\right)$. From (2.1), we have an exact sequence
@@
which implies that $\mathfrak{p}R_{\mathfrak{p}}\in\operatorname{Ass}_{R_{\mathfrak{p}}}(\mathcal{T}_{k}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))$ and hence
@@
Since $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$ for all $11\leq i\leq k-$ 1, we see from [52, 5.6] that
@@
We claim that $k\leq 0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$. Indeed, if $k>\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$, then by assumption $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))<\infty$, and from the Auslander-Bridger formula, $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))<k$. But since $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$ for all $11\leq i\leq k-$ 1, we get that $\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$ is a totally reflexive. This shows that $\operatorname{Ext}_{R_{\mathfrak{p}}}^{k}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$, which does not occurs.
@@
Observe that Theorem 3.2 is the version with the dual of an $R$-module being of finite Gorenstein dimension of [14, Proposition 2.4] when $C=R$. Now, we explore some consequences of Theorem 3.2.
@@
Proof.
The equivalences $(1)\Leftrightarrow(2)\Leftrightarrow(3)$ follow from Theorem 3.2. The equivalence $(2)\Leftrightarrow(4)$ follows from (2.1). ∎
This theorem allows us to recover one of the results described in the main result of [52].
3.5 Corollary ([52, Theorem 5.8(1)]).
Let $R$ be a local ring, $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $k$ be a non-negative integer. If $M$ satisfies $(\widetilde{S}_{k})$, then $M$ is $k$-torsionfree.
Note that Theorem 3.2 provides an answer to the following question posed in [37].
3.6 Question ([37, Question 1.1]).
@@
The following result is an analogous version of [44, Theorem 4.5], which is one of the main results of that paper.
@@
    $M_{\mathfrak{p}}$ is $(k-1)$-torsionfree for all $\mathfrak{p}\in\widetilde{X}^{k-1}(R)$.
@@
Reciprocally, suppose that (1) and (2) hold. By Theorem 3.2, it is sufficient to show that $M$ satisfies $(\widetilde{S}_{k}),$ that is $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$. By (2), it holds if $\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\geq k$. Now, let $10pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\leq k-$ 1. Then by the Auslander-Bridger formula, $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))\leq k-1$. Thus, by (1), $\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$ is totally reflexive, and hence $M_{\mathfrak{p}}$ as well. In particular, $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})=\min\{k,\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$. ∎
@@ 4. Characterization of totally reflexives
In this section, $R$ is a local ring. Next, we derive some criteria of totally reflexive modules, assuming the condition $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, and involving the vanishing of the Ext functor. We first provide a general criterion via the vanishing of $\operatorname{Ext}_{R}^{1\leq i\leq\operatorname{depth}_{R}(R)}(M,R)$, and subsequently, we provide criteria over Cohen-Macaulay local rings and later for locally totally reflexive modules on the punctured spectrum.
@@
$(1)\Rightarrow(2)$. According to Corollary 3.5, $M$ is $t$-torsion-free, meaning that $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $1\leq i\leq t$. Since $\operatorname{G-dim}_{R}(\operatorname{Tr}(M))\leq t,$ it follows that $\operatorname{Tr}(M)$ is totally reflexive. Hence, $M$ is totally reflexive.
$(2)\Rightarrow(3)$. This follows from definition of totally reflexives modules.
$(3)\Rightarrow(1)$. Suppose that $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t$. Since $M\approx\operatorname{Tr}(\operatorname{Tr}(M))$, this implies that $\operatorname{Tr}(M)$ is $t$-torsion-free. Given that $\operatorname{G-dim}_{R}(\operatorname{Tr}(M))<\infty$, according to [50, Theorem 42], $\operatorname{Tr}(M)$ satisfies $(\widetilde{S}_{t})$. Thus, $\operatorname{depth}_{R}(\operatorname{Tr}_{R}(M))\geq t$ since $0pt_{R}(R)=t$. By the Auslander-Bridger formula, $\operatorname{Tr}(M)$ is totally reflexive. Thus, $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}_{R}(M),R)=0$ for all $i>0$. In particular, $M$ is $t$-torsionfree, and by [50, Proposition 11], $M$ satisfies $(\widetilde{S}_{t})$. ∎
@@
As an immediate application of Theorem 4.1, we obtain the following characterization for a module over a Cohen-Macaulay ring to be totally reflexive, involving the vanishing of the Ext functor.
4.2 Theorem.
Let $R$ be a Cohen-Macaulay local ring of dimension $d$ and let $M$ be a non-zero $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following are equivalents:
@@
Proof.
Since $R$ is Cohen-Macaulay, it is clear that $M$ is maximal Cohen-Macaulay if and only if $M$ satisfies $(\widetilde{S}_{d})$. Thus, the result follows immediately from Theorem 4.1. ∎
@@
It is known that if $R$ is a Gorenstein local ring and $M$ is maximal Cohen-Macaulay, then $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i>0$. We can observe from Theorem 4.2, that in this fact instead of $R$ being Gorenstein, we can consider the weaker condition of $\operatorname{G-dim}_{R}(M^{\ast})<\infty$.
@@
Let $R$ be a Cohen-Macaulay local ring and let $M$ be a Cohen-Macaulay $R$-module such that $M^{\ast}\not=0$ and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then $M$ is totally reflexive.
@@
Since $M^{\ast}\not=0$, note that $\operatorname{grade}_{R}(M)=0$. Moreover, by [7, Corollary 2.1.4], $\operatorname{grade}_{R}(M)=\operatorname{dim}_{R}(R)-\operatorname{dim}_{R}(M)$ and hence $\operatorname{dim}_{R}(M)=\operatorname{dim}_{R}(R).$ Thus from Cohen-Macaulayness of $M$, we see that $M$ is maximal Cohen-Macaulay. Therefore, Theorem 4.2 asserts that $M$ is totally reflexive. ∎
Now, as an application of Theorem 4.2, we extend the formula presented in [13, Lemma 4.1] to all $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. But before, we recall the definition of reduced grade. The reduced grade of an $R$-module $M$ is defined as
@@
Let $R$ be a ring. If $\operatorname{grade}_{R}(M)>0,$ then $\operatorname{grade}_{R}(M)=\operatorname{r.grade}_{R}(M).$
4.6 Theorem.
@@
    $M$ is not maximal Cohen-Macaulay and $\operatorname{dim}(M)=\operatorname{dim}(R)$.
@@
    $M$ is not maximal Cohen-Macaulay and $\operatorname{dim}(M)\not=\operatorname{dim}(R)$.
@@
The equality in the case (1) was proved in [13, Lemma 4.1].
Now, consider the case (2). Since $\operatorname{dim}(M)\not=\dim(R)$, we see from [7, Theorem 3.5.7] that $\sup\{i\geq 0:H_{\mathfrak{m}}^{i}(M)\not=0\text{ and }i\not=d\}=\operatorname{dim}(M).$ Moreover, by [7, Corollary 2.14], we get that $\operatorname{grade}_{R}(M)=d-\operatorname{dim}_{R}(M)>0$. Then by Remark 4.5, $\operatorname{r.grade}_{R}(M)=d-\operatorname{dim}(M)$. Now observe that the desired equality holds.
Finally, let us prove the equality in the case (3). In this case, by [7, Theorem 3.5.7], $\sup\{i\geq 0:H_{\mathfrak{m}}^{i}(M)\not=0\text{ and }i\not=d\}=-\infty$. On the other hand, by Theorem 4.2, $M$ is totally reflexive and hence $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i\geq 1$. Thus $\operatorname{r.grade}_{R}(M)=\infty$ and we can see that the equality holds. ∎
@@
We may assume that $R$ is complete. Because of [7, Theorem 3.5.8], the desired equality is a reformulation of the equality (4.1). ∎
@@
Let $R$ be a Cohen-Macaulay local ring of dimension $d$ and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Let $n$ be a non-negative integer such that $\operatorname{depth}_{R}(M)\geq n$. If $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d-n,$ then $M$ is totally reflexive.
@@
Since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d-n$ and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, by Corollary 4.7, $\operatorname{Ext}_{R}^{i}(M,\omega)=0$ for all $1\leq i\leq d-n$. Since [7, Exercise 3.1.24] say us that
@@
then $d-\operatorname{depth}_{R}(M)=0$ or $d-\operatorname{depth}_{R}(M)>d-n$. The second case does not occur because $\operatorname{depth}_{R}(M)\geq n$ by assumption. Thus $d=\operatorname{depth}_{R}(M)$ and hence $M$ is maximal Cohen-Macaulay. Thus, by Theorem 4.2, $M$ is totally reflexive.
@@
The following result provides a reflexivity criterion for $R$-modules $M$ with $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ of locally totally reflexives on the punctured spectrum of $R$.
@@
Let $R$ be a local ring of depth $t$ and let $M$ be a non-zero $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Suppose that $M$ is locally totally reflexive on the punctured spectrum of $R$. Then $0pt_{R}(M)\leq t$ with equality if and only if $M$ is totally reflexive.
@@
First we prove that $\operatorname{depth}_{R}(M)\leq t$. Set $r=\operatorname{depth}_{R}(M)$ and suppose that $r>t$. Then $M$ satisfies $(\widetilde{S}_{r})$ since $M$ is locally totally reflexive on the punctured spectrum of $R$. In particular, $M$ satisfies $(\widetilde{S}_{t})$. By Theorem 4.1, $M$ is totally reflexive, which contradicts the fact of $r>t$.

### v5

@@ 3. A generalization of a theorem of Auslander and Bridger
Let $R$ be a local ring of depth $t$, $M$ be a non-zero $R$-module and $0\leq k\leq t$ be an integer. If $M$ is a $k$-syzygy of an $R$-module of depth 0, then $\operatorname{depth}_{R}(M)=k$.
@@
It is sufficient to consider $k\geq 1$. We proceed by induction on $k$. Suppose $k=1$. Then $t\geq 1$ and there exists an exact sequence $0\to M\to F_{0}\to L\to 0,$ where $F_{0}$ is free and $L$ has depth 0. The depth lemma gives us the inequalities
@@
Since $t\geq 1$, it follows from (3.1) that $\operatorname{depth}_{R}(M)\leq 1$. Thus (3.2) yields that $\operatorname{depth}_{R}(M)=1$.
Now suppose $1<k\leq t$. From the given hypotheses, there exists an exact sequence
@@
where each $F_{i}$ is free and $\operatorname{depth}_{R}(L)=0$. Then it induces exact sequences
@@
(3.4) $0\to N\to F_{k-2}\to\cdots\to F_{0}\to L\to 0.$
If $N=0$, then (3.4) says that $\operatorname{pd}_{R}(L)\leq k-2$ and, by the Auslander-Buchsbaum formula, $t\leq k-2,$ which does not occurs by assumption. So, $N\not=0$. Then by induction, we have that $\operatorname{depth}_{R}(N)=k-1$. Applying the depth lemma in (3.3) we get that
@@
Thus, if $t\leq\operatorname{depth}_{R}(M)-1$, then $k-1\geq t$, which contradicts our assumption. Therefore, we must have $t>\operatorname{depth}_{R}(M)-1$. Consequently, $\operatorname{depth}_{R}(M)-1\leq k-1$, which implies $\operatorname{depth}_{R}(M)\leq k$. Therefore, $k=\operatorname{depth}_{R}(M)$. ∎
Below, we provide a characterization of a module to be $k$-torsionfree or $k$-syzygy, involving the condition $(\widetilde{S}_{k})$. This characterization generalizes [Syzygies, Theorem 3.8]
@@
Let $R$ be a ring and let $k$ be a non-negative integer. Let $M$ be an $R$-module such that $M^{\ast}$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(R)$. Then the following conditions are equivalent:
@@
The implications $(1)\Rightarrow(2)\Rightarrow(3)$ follow from [GorensteinDimensionAndTorsionOfModulesOverCommutativeNoetherianRings, Proposition 38]. Now let us prove $(3)\Rightarrow(1)$. We prove this by induction on $k$. If $k=0$, there is nothing to prove. Assume that $k\geq 1$. By induction, $M$ is (k1)(k-1)-torsionfree, that is $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $1\leq i\leq k-1$, and so it remains to conclude that $\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R)=0$.
By contradiction, assume that $\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R)\not=0$. Let $\mathfrak{p}\in\operatorname{Ass}_{R}(\operatorname{Ext}_{R}^{k}(\operatorname{Tr}(M),R))$. Then $\mathfrak{p}R_{\mathfrak{p}}\in\operatorname{Ass}_{R_{\mathfrak{p}}}\left(\operatorname{Ext}_{R_{\mathfrak{p}}}^{k}\left(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}}\right)\right)$. For a Noetherian local ring $S$ and a (finitely generated) $S$-module NN, the $S$-module $\operatorname{Ext}_{S}^{1}(N,S)$ is a submodule of $\operatorname{Tr}(N)$. Thus, taking $S=R_{\mathfrak{p}}$ and $N=\Omega_{R_{\mathfrak{p}}}^{k-1}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))$, we obtain an exact sequence
@@
which implies that $\mathfrak{p}R_{\mathfrak{p}}\in\operatorname{Ass}_{R_{\mathfrak{p}}}(\mathcal{T}_{k}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))$, and hence
@@
Since $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$ for all $1\leq i\leq k-1$, we see from [HomologicalDimensionsOfRigidModules, 5.6] that
@@
We claim that $k\leq 0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$. Indeed, if $k>\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$, then by assumption $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))<\infty$, and by the Auslander-Bridger formula, $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))<k$. But since $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$ for all $1\leq i\leq k-1$, we get that $\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$ is a totally reflexive. This shows that $\operatorname{Ext}_{R_{\mathfrak{p}}}^{k}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}),R_{\mathfrak{p}})=0$, which does not occurs.
@@
Observe that Theorem 3.2 is the version with the dual of an $R$-module being of finite Gorenstein dimension of [LinkageOfModulesAndTheSerreConditions, Proposition 2.4] when $C=R$. Now, we explore some consequences of Theorem 3.2.
@@
As another consequence, this theorem allows us to recover [HomologicalDimensionsOfRigidModules, Theorem 5.8(1)]. Additionally, it provides an answer to the following question posed in [WhenArenSysymodulesnTorsionfree].
3.5 Question ([WhenArenSysymodulesnTorsionfree, Question 1.1]).
@@
The following result is an analogous version of [ktorsionlessmoduleswithfiniteGorensteindimension, Theorem 4.5], which is one of the main results of that paper.
@@
    $M_{\mathfrak{p}}$ is (k1)(k-1)-torsionfree for all $\mathfrak{p}\in\widetilde{X}^{k-1}(R)$.
@@
Reciprocally, suppose that (1) and (2) hold. By Theorem 3.2, it is sufficient to show that $M$ satisfies $(\widetilde{S}_{k}),$ that is $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$. Let $\mathfrak{p}\in\operatorname{Supp}(M)$. By (2), it holds if $\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\geq k$. Now, let $0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\leq k-1$. Then by the Auslander-Bridger formula, $\operatorname{G-dim}_{R_{\mathfrak{p}}}(\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}))\leq k-1$. Thus, by (1), $\operatorname{Tr}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$ is totally reflexive, and hence $M_{\mathfrak{p}}$ as well. In particular, $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})=\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})=\min\{k,\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$. ∎
@@ 4. Characterization of totally reflexives
In this section, $R$ is a local ring. Next, we derive some criteria for an $R$-module $M$ to be totally reflexive, assuming the condition $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and involving the vanishing of the Ext functor. We first provide a general criterion via the vanishing of $\operatorname{Ext}_{R}^{1\leq i\leq\operatorname{depth}_{R}(R)}(M,R)$, and subsequently, we provide criteria over Cohen-Macaulay local rings and later for locally totally reflexive modules on the punctured spectrum.
@@
$(1)\Rightarrow(2)$. According to Theorem 3.2 or [HomologicalDimensionsOfRigidModules, Theorem 5.8(1)], $M$ is $t$-torsionfree, meaning that $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $1\leq i\leq t$. Since $\operatorname{G-dim}_{R}(\operatorname{Tr}(M))\leq t,$ it follows that $\operatorname{Tr}(M)$ is totally reflexive. Hence, $M$ is totally reflexive.
$(2)\Rightarrow(3)$. This follows from the definition of totally reflexive modules.
$(3)\Rightarrow(1)$. Suppose that $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t$. Since $M\approx\operatorname{Tr}(\operatorname{Tr}(M))$, this implies that $\operatorname{Tr}(M)$ is $t$-torsionfree. Given that $\operatorname{G-dim}_{R}(\operatorname{Tr}(M))<\infty$, according to [GorensteinDimensionAndTorsionOfModulesOverCommutativeNoetherianRings, Theorem 42], $\operatorname{Tr}(M)$ satisfies $(\widetilde{S}_{t})$. Thus, $\operatorname{depth}_{R}(\operatorname{Tr}(M))\geq t$ since $0pt_{R}(R)=t$. By the Auslander-Bridger formula, $\operatorname{Tr}(M)$ is totally reflexive. Thus, $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $i>0$. In particular, $M$ is $t$-torsionfree, and by [GorensteinDimensionAndTorsionOfModulesOverCommutativeNoetherianRings, Proposition 11], $M$ satisfies $(\widetilde{S}_{t})$. ∎
@@
As an immediate application of Theorem 4.1, we obtain the following characterization for a module over a Cohen-Macaulay local ring to be totally reflexive, involving the vanishing of the Ext functor.
4.2 Corollary.
Let $R$ be a Cohen-Macaulay local ring of dimension $d$, and let $M$ be a non-zero $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following are equivalent:
@@
It is known that if $R$ is a Gorenstein local ring and $M$ is maximal Cohen-Macaulay, then $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i>0$. We can observe from Corollary 4.2, that in this fact instead of $R$ being Gorenstein, we can consider the weaker condition of $\operatorname{G-dim}_{R}(M^{\ast})<\infty$.
@@
Let $R$ be a Cohen-Macaulay local ring, and let $M$ be a Cohen-Macaulay $R$-module such that $M^{\ast}\not=0$ and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then $M$ is totally reflexive.
@@
Since $M^{\ast}\not=0$, note that $\operatorname{grade}_{R}(M)=0$. Moreover, by [bruns, Corollary 2.1.4], $\operatorname{grade}_{R}(M)=\operatorname{dim}(R)-\operatorname{dim}_{R}(M)$ and hence $\operatorname{dim}_{R}(M)=\operatorname{dim}(R).$ Thus from Cohen-Macaulayness of $M$, we see that $M$ is maximal Cohen-Macaulay. Therefore, Corollary 4.2 asserts that $M$ is totally reflexive. ∎
Now, as an application of Corollary 4.2, we extend the formula presented in [LinkageOfFiniteGorensteinDimensionModules, Lemma 4.1] to all $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. But before, we recall the definition of reduced grade. The reduced grade of an $R$-module $M$ is defined as:
@@
Let $R$ be a ring and let $M$ be an $R$-module. If $\operatorname{grade}_{R}(M)>0,$ then $\operatorname{grade}_{R}(M)=\operatorname{r.grade}_{R}(M).$
4.6 Proposition.
@@
    $M$ is not maximal Cohen-Macaulay and $\operatorname{dim}_{R}(M)=\operatorname{dim}(R)$.
@@
    $M$ is not maximal Cohen-Macaulay and $\operatorname{dim}_{R}(M)\not=\operatorname{dim}(R)$.
@@
The equality in the case (1) was proved in [LinkageOfFiniteGorensteinDimensionModules, Lemma 4.1].
Now, consider the case (2). Since $\operatorname{dim}_{R}(M)\not=\dim(R)$, we see from [bruns, Theorem 3.5.7] that $\sup\{i\geq 0:H_{\mathfrak{m}}^{i}(M)\not=0\text{ and }i\not=d\}=\operatorname{dim}_{R}(M).$ Moreover, by [bruns, Corollary 2.1.4], we get that $\operatorname{grade}_{R}(M)=d-\operatorname{dim}_{R}(M)>0$. Then by Remark 4.5, $\operatorname{r.grade}_{R}(M)=d-\operatorname{dim}_{R}(M)$. Now observe that the desired equality holds.
Finally, let us prove the equality in the case (3). In this case, by [bruns, Theorem 3.5.7], $\sup\{i\geq 0:H_{\mathfrak{m}}^{i}(M)\not=0\text{ and }i\not=d\}=$-$\infty$. On the other hand, by Corollary 4.2, $M$ is totally reflexive and hence $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i\geq 1$. Thus $\operatorname{r.grade}_{R}(M)=\infty$ and we can see that the required equality holds. ∎
@@
We may assume that $R$ is complete. Because of [bruns, Theorem 3.5.8], the desired equality is a reformulation of the equality (4.1). ∎
@@
Let $R$ be a Cohen-Macaulay local ring of dimension $d$, and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Let $n$ be a non-negative integer such that $\operatorname{depth}_{R}(M)\geq n$. If $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d-n,$ then $M$ is totally reflexive.
@@
Since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d-n$ and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, by Corollary 4.7, $\operatorname{Ext}_{R}^{i}(M,\omega)=0$ for all $1\leq i\leq d-n$. Since [bruns, Exercise 3.1.24] says us that
@@
then $d-\operatorname{depth}_{R}(M)=0$ or $d-\operatorname{depth}_{R}(M)>d-n$. The second case does not occur because $\operatorname{depth}_{R}(M)\geq n$ by assumption. Thus $d=\operatorname{depth}_{R}(M)$ and hence $M$ is maximal Cohen-Macaulay. Thus, by Corollary 4.2, $M$ is totally reflexive.
@@
The following result provides a totally reflexivity criterion for $R$-modules $M$ with $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ that are locally totally reflexives on the punctured spectrum of $R$.
@@
Let $R$ be a local ring of depth $t$, and let $M$ be a non-zero $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Suppose that $M$ is locally totally reflexive on the punctured spectrum of $R$. Then $0pt_{R}(M)\leq t$ with equality if and only if $M$ is totally reflexive.
@@
First, we prove that $\operatorname{depth}_{R}(M)\leq t$. Set $r=\operatorname{depth}_{R}(M)$ and suppose that $r>t$. Then $M$ satisfies $(\widetilde{S}_{r})$ since $M$ is locally totally reflexive on the punctured spectrum of $R$. In particular, $M$ satisfies $(\widetilde{S}_{t})$. By Theorem 4.1, $M$ is totally reflexive, which contradicts the fact that $r>t$.